Logbooks Lab Maintenance Evaporator_1 Evaporator_2 Laser cutter Target Production Test-Stand RH-ISAC RH-Cyclotron RH-Meson Hall RH-Beamlines RH-ARIEL
  RH-ISAC, Page 94 of 140  Not logged in ELOG logo
ID Date Author Category Type Module Target/Number Subject Text Attachments
  927   Tuesday, June 02, 2015, 14:32 David WangSouth Hot-CellStandard OperationTM4SiC #31Leak check on TM4 HS water line with new SiC #31 target. at SHCChad and I leak checked TM4 HS water line
with SiC #31 target installed. leak rate:0.0xE-9atm.cc/sec. 
No leak.  The line will be helium pressurized
 003.JPG005.JPG001.JPG 
  926   Tuesday, June 02, 2015, 11:54 Travis CaveConditioning StationStandard OperationTM4SiC#31Module movedTM#4 with SiC#31 has been moved from the
south hot cell to the conditioning station.
  
  925   Tuesday, June 02, 2015, 10:39 Travis CaveSouth Hot-CellStandard OperationTM4SiC#31Electrical CheckElectrical check on TM#4 with SiC#31. See
attached PDF for details.
 TM4_SiC31_elecchk_SHC.PDF 
  924   Tuesday, June 02, 2015, 09:43 chad fisherSouth Hot-CellStandard OperationTM4SiC#31SiC#31 installedTa #45 has been removed from TM4 and SIC#31installed.   
  923   Tuesday, June 02, 2015, 06:51 David WangSouth Hot-CellStandard OperationTM4 Target Hall Scheduled Activities April 07 - June 26 2015Target Hall schedule weekly update. All
jobs are finished on time so far.
 Target_Hall_Scheduled_Activities__Apr_07_to_June_26_2015.mppTarget_Hall_Scheduled_Activities__Apr_07_to_June_26_2015.pdf 
  922   Monday, June 01, 2015, 14:24 Travis CaveSouth Hot-CellStandard OperationTM4Ta#45Electrical CheckPost beam electrical check on the Ta#45
target installed on TM#4. See attached PDF.
 TM4_Ta45_elecchk_postbeam.PDF 
  921   Monday, June 01, 2015, 11:23 Travis CaveITEStandard OperationTM4Ta#45Module movedTM#4 with Ta#45 has been moved from the
east target station to the south hot cell.
  
  920   Friday, May 29, 2015, 09:16 Travis CaveConditioning StationStandard OperationTM2no targetModule movedTM#2 with no source tray has been moved
from the conditioning station to the north
east silo.
  
  919   Wednesday, May 27, 2015, 08:57 Travis CaveConditioning StationStandard OperationTM2 Module movedTM#2 with no source tray has been moved
from the south hot cell to the conditioning
station.
  
  918   Tuesday, May 26, 2015, 14:34 chad fisherSouth Hot-CellRepairTM2 Shutter replacementThe shutter in TM2 was successfully replaced
today.
New lock washer and nut were installed
 Shutter.jpg 
  917   Tuesday, May 26, 2015, 11:49 Grant MinorAssembly/Bench TestRepairTM2 TM2 New Source Tray - Braze & Solder Report - 22 May 2015Hello,

For reference, I have attached a list
 6x 
  916   Friday, May 22, 2015, 15:44 David WangITWStandard OperationTM1UCx125 Shielding blocks have been transferred from ITE to ITW.5 shielding blocks have been transferred
from ITE to ITW. I  checked all ITW
water signals, and they are good.  ITW
  
  915   Wednesday, May 13, 2015, 13:55 chad fisherSouth Hot-CellRepairTM2 Shutter removalTM2 has been successfully de-shuttered.
All went pretty much according to
plan except I had to remove one fastener
 TM2_20150513_131339.jpgTM2_20150513_134232.jpg 
  914   Monday, May 11, 2015, 11:20 David WangSouth Hot-CellStandard OperationTM2 Target Hall Scheduled Activities April 07 - June 26 2015The schedule has some changes due to 
new TM2 source tray coil cooling line leaking
at test stand. Now,It is re-scheduled in
 6x 
  913   Monday, May 11, 2015, 10:45 Anders MjosConditioning StationMaintenance  RGA electronics box installed The communications chip has been
replaced. The electronics box has been grounded
(see attached pdf).
 2015-05_RGA_GND_connection.pdf 
  912   Thursday, May 07, 2015, 11:02 Travis CaveSouth Hot-CellRepairTM2No targetWiring of TM#2Wires in TM#2 have been terminated at the
top of the module.
  
  911   Wednesday, May 06, 2015, 14:30 Travis CaveSouth Hot-CellRepairTM2No targetWiring of TM#2Wires for TM#2 steering elements, and png
gauge have been run down to the containment
box level. Chad has placed them in the little
 Photo_2015-05-06_8_26_10_AM.JPGPhoto_2015-05-06_8_26_17_AM.JPG 
  910   Tuesday, May 05, 2015, 10:57 chad fisherSouth Hot-CellRepairTM2 TM2 WiringOld wiring was removed yesterday, new wiring
run thru the vacuum chase. Old IMG gauge
removed today, new cable trays installed
 TM2_New_Wiring_20150504_13.jpgTM2_Wiring_20150505_10.jpg 
  909   Monday, May 04, 2015, 13:32 Anders MjosConditioning StationMaintenance  TK2 expansion tank topped upCaled cyclotron ops to open P23 and confirmed
penthouse valve open.
Filled up TK2 to 21 in water leve.
  
  908   Friday, May 01, 2015, 13:48 David WangSouth Hot-CellStandard OperationTM2 Target Hall Scheduled Activities April 07 - June 26 2015 updateTarget Hall Scheduled Activities 
April 07 - June 26  2015 update.  
   2015-05-01. All TH scheduled
 Target_Hall_Scheduled_Activities__Apr_07_to_June_26_2015.mppTarget_Hall_Scheduled_Activities__Apr_07_to_June_26_2015.pdf 
ELOG V2.9.2-2455